1. Given a set of numbers: A = {1, 3, 5, 2, 8, 7, 9, 4 }.

Add each digit of your N-Number (e.g. N0 0 3 2 2 0 1 8) to each element of A

separately to get your own final array B. ( e.g. B={1,3,8,4,10,7,10,12} = {B1,B2,B3,B4,B5,B6,B7,B8}. )

a) Your set B = ?

b) Use merge-sort to sort the array B. Show each step.

c) How many “comparisons” for each algorithm above?

2. Use binary-search for finding the number 6 for the sorted array based on Question 1, show each step. (6 might not be in your sorted array.)

3. The time complexity equation of merger-sort is T(n) = 2* T(n/2) + n, where T(1) = C and C is a constant. Solve this equation by giving detailed steps. (Hint: See lecture notes).

4. How do we use linked-lists to represent adjacency-lists in a graph? Provide an example including at least 6 vertices and 8 edges.
Answer #1

1. A:1 3.5.2.8.7.9.0} | lat Us {o.1,21,0,0,1,2} | a) | 。 - {1,0,73,8 LL, 10,6 [urs all lol ill Lo 6 01 3 lo 白白向自西向立 11 6 ( 3c

1. 2. In sorting a shows the number of compare- Son at that step. val no. of comparisons - Ild+1+1/3+370 - 15 Array B= {1, 3

after k function calls, fin నందం T(n) = 2 + ) + (n +n un to---k times) @ T(n) = 2*T (2/2) + kn - for 2 / 2 o ka logan By subs

IS A rayfoorsday + - Rorovo/pro H ISS FIT 1 / Ft FTT FIF Is | 기 FTFITF 7) Ft F | [DIY ETFT- [ ELES FTS E To | |

If you're still having any doubt then please feel free to ask in the comment section.